OpenStudy (anonymous):

Assume that n(U) = 90, n(A) = 40, and n(B) = 45. Then n( A′∪B) = ??

OpenStudy (anonymous):

I thought A' would be 50 because 90-40=50 and B is 45, so 45+50=95, but apparently that isn't correct!

OpenStudy (phi):

U is the universal set. It is everything. So you can't get more than 90 if you get 95, it must mean that some of B are also in A.

OpenStudy (anonymous):

A^c = (U-A)

OpenStudy (anonymous):

Clearly I wasn't thinking! Alright, so how would I find the answer for this?

OpenStudy (shaik0124):

90-(40+45)=90-85=5

OpenStudy (anonymous):

5 would be what in this?

OpenStudy (shaik0124):

answer of n(a' union b)

OpenStudy (anonymous):

I think I get it. Thanks!

OpenStudy (shaik0124):

@Wislar did u got it

OpenStudy (anonymous):

This doesn't seem to be working out. Wouldn't 5 actually be (A U B)'?

OpenStudy (anonymous):

Would it be 50?

OpenStudy (phi):

Is there any more info? Because this picture (for example) seems to fit:  which gives 80 as the answer